### v1

# Robust and Smooth Estimation of the Extreme Tail Index via Weighted Minimum Density Power Divergence Saida Mancer, Abdelhakim Necir Djamel

Robust and Smooth Estimation of the Extreme Tail Index via Weighted
Minimum Density Power Divergence
Saida Mancer, Abdelhakim Necir Djamel
Meraghni
Laboratory of Applied Mathematics, Mohamed Khider University, Biskra, Algeria
By introducing a weight function into the density power divergence, we develop a new class of robust and smooth estimators for the tail index of Pareto-type distributions, offering improved efficiency in the presence of outliers. These estimators can be viewed as a robust generalization of both weighted least squares and kernel-based tail index estimators. We establish the consistency and asymptotic normality of the proposed class. A simulation study is conducted to assess their finite-sample performance in comparison with existing methods.
Keywords: Extreme value index; Least squares estimation; Minimum density power divergence; Robustness.
AMC 2020 Subject Classification: Primary 62G32; 62G05; 62G20; 62G35.
Corresponding author:
ah.necir@univ-biskra.dz
E-mail address:
mancer.saida731@gmail.com (S. Mancer)
djamel.meraghni@univ-biskra.dz (D. Meraghni)
1. Introduction
Let be independent and identically distributed (iid) of non-negative random variables (rv’s) viewed as copies of a rv defined over some probability space with cumulative distribution function (cdf) We assume that the tail distribution is regularly varying at infinity with negative index i.e., for every
The parameter often referred to as the shape parameter, tail index, or extreme value index (EVI), plays a fundamental role in extreme value analysis, as it characterizes the heaviness of the distribution’s right tail. The estimation of has received considerable attention over the past four decades. The most widely used estimator of is Hill’s estimator (Hill, 1975) defined as
where denote the order statistics pertaining to the sample and is an integer sequence satisfying and as The discrete nature and lack of stability of Hill’s estimator constitute significant drawbacks. In particular, the inclusion of a single additional upper-order statistic, i.e. increasing by one, can lead to substantial deviations from the true value of the tail index, reflecting the estimator’s sensitivity to the choice of the threshold As a result, plotting this estimator as a function of the number of upper-order statistics typically produces a zigzag pattern, which leads to instability and fluctuations in determining the optimal sample fraction used in the estimation of the computation of To address this issue, Csörgő et al. (1985) (CDM) introduced more general weights in place of the natural weight appearing in the second expression of Hill’s estimator leading to the definition of the following kernel-type estimator:
where is a continuous nonnegative nonincreasing function on such that For their part, Hüsler et al. (2006) used the weighted least squares estimator (WLSE) given by
where and is a suitable continuous nonnegative nonincreasing function defined on and vanishes elsewhere such that Similar estimators to are also considered in Beirlant et al. (1996), Viharos (1999) and recently Caeiro et al. (2021). The authors pointed out that the least squares estimator may be rewritten into CDM’s one for the kernel function
For example, by choosing we obtain
However, the converse does not necessarily hold; that is may be a kernel function without being one. For instance, if we take then the corresponding function is identically zero. Here, denotes the indicator function of the set The most commonly used weighted functions the indicator function as well as the Epanechnikov, biweight, triweigh and quadweight weighted functions which are defined on the interval by
and zero elsewhere respectively. For the use of this type of weight functions, we refer the reader to Groeneboom et al. (2003) and Hüsler et al. (2006).
The widely adopted approach for estimating the parameters of an extreme value distribution in extreme value analysis is the maximum likelihood estimation (MLE) method. While MLEs enjoy desirable asymptotic properties, they can be highly sensitive to outliers or deviations from the assumed extreme value models Brazauskas and Serfling (2000). Robust statistics offer a valuable alternative for mitigating the influence of outliers, and consequently, reducing deviations from the underlying parametric models. Incorporating robust statistical methods into extreme value theory has been shown to enhance both the accuracy and reliability of parameter estimates. (Dell’Aquila and Embrechts, 2006). Juárèz and Schucany (2004) appear to be the first to employ the Minimum Density Power Divergence (MDPD) introduced by Basu et al. (1998) for the robust estimation of parameters in extreme value distributions. Since then, this divergence measure has become one of the most widely used tools for robust parameter estimation in the context of extreme value theory. Several authors, including Kim and Lee (2008), Dierchx et al. (2013), Goegebeur et al. (2014), Dierckx et al. (2021) have employed the MDPD framework to estimate the tail index and quantiles of Pareto-type distributions.
Recently, Ghosh (2017) proposed a robust MDPD-based estimator for a real-valued tail index, serving as a robust generalization of the estimator introduced by Mattys and Beirlant (2003). This approach also accounts for the non-identically distributed structure of the exponential regression model, following the methodology of Ghosh and Basu (2013). Additionally, Dierchx et al. (2013) applied the MDPD framework to an extended Pareto distribution to model relative excesses over a high threshold. More recently, Minkah el al. (2023) developed a robust estimator for the tail index of a Pareto-type distribution by applying the MDPD approach within the exponential regression model framework.
We are thus faced with two main challenges: robustness and the instability of the optimal sample fraction used in the computation of tail index estimators. To address these issues, we propose assigning a weight function to the density power divergence, leading to a robust generalization of the weighted least squares tail index estimator To the best of our knowledge, this approach has not yet been explored in the existing literature. The proposed estimation procedure is presented in detail in Section
The remainder of this paper is organized as follows. Section 2 provides a brief overview of the MDPD estimation method, originally introduced by Basu et al. (1998). The proposed weighted MDPD-based estimation of the tail index is developed in Section 3. Section 4 presents the asymptotic properties of the proposed estimator, specifically its consistency and asymptotic normality, with detailed proofs deferred to Section 6. The finite-sample performance of the estimator is evaluated in Section 5 through a simulation study, including comparisons with existing estimators. Section 8 collects several technical lemmas and propositions, while Section 9 gathers the figures corresponding to the simulation results.
2. Minimum density power divergence
Given two probability densities and Basu et al. (1998) introduced a new distance between them called the density power divergence
where is a nonnegative tuning parameter. The case corresponding to is obtained from the general case by letting leading to the classical Kullback-Leibler divergence denoted Let us consider a parametric model of densities and suppose that we consider the estimation of the parameter Let be the cdf corresponding to the density The minimum density power divergence (MDPD) is a functional defined by It is clear that the term in does not contribute in the minimization of over Then minimization in the computation of the MDPD functional reduces to
Given a random sample from the distribution we may estimate the objective function in by substituting with its empirical counterpart For a given tuning parameter the MDPD estimator of may be obtained by minimizing (over the quantity
which in turn equals
Thus the MDPD estimator of minimizing will be a solution of the following equation
The tuning parameter plays a crucial role in the MDPD framework, as it governs the trade-off between efficiency and robustness. Specifically, when is close to zero, the estimator is more efficient but less robust to outliers. Conversely, as increases, robustness improves at the cost of some loss in efficiency. It has been shown that estimators with small values of can exhibit strong robustness properties while incurring only a slight loss in asymptotic efficiency relative to the maximum likelihood estimator under correct model specification.
3. Weighted MDPD estimation of the tail index
As stated in Lemma we have
as where
and is the weight function introduced in From equation we deduce that for large represents the tail functional associated with the tail index estimator defined in Since vanishes at and is continuous, it is more appropriate to use instead of where
Accordingly, the integral in becomes
This modification will be adopted throughout the empirical estimation of tail functionals. Specifically, in we replace the threshold with the intermediate order statistic and the cdf with its empirical counterpart
to obtain
which corresponds exactly to the expression of On the other hand, under appropriate assumptions on the weight function we stated in Lemma 8.1 that assertion is equivalent to
as or equivalently
as where
and
denotes the standard Pareto density function. From this, we infer that the integral equation above leads to the same tail functional In other words, the estimator may be viewed as the solution to the empirical counterpart of that equation, or equivalently which reads:
In conclusion, the tail index estimator can be viewed as the result of minimizing the Kullback–Leibler divergence between two weighted densities, and This leads us to consider the general case of the weighted minimum density power divergence (WMDPD), between and by
in the sense that
Since the term in plays no role in the minimization of over it is sufficient to minimize the simplified divergence functional
for sufficiently large In the other terms, the WMDPD of minimizing is solution of the following equation
By substituting with and with in the previous equation, we en up with the WMDPD estimator of tail index as solution of the equation
We may therefore regard for as a robust generalization of It is also worth noting that can be seen as a weighted extension of the MDPD estimator previously introduced by Dierchx et al. (2013).
4. Main results
Next, we establish the existence and consistency of a sequence of solutions to the estimating equation for To this end, we consider a class of weight functions that satisfy the following regularity conditions:
is continuous nonincreasing nonnegative on the interval with
and their -th derivatives are bounded on the interval
as well as their -th derivatives are bounded on the interval
The regularity assumptions are relatively mild, as the commonly used weight functions listed satisfy them. In fact, Table 1 presents the explicit expressions of the function and its first derivative for each of the aforementioned weight functions. One can verify that both and are bounded over the interval and that their second and third derivatives are also bounded. As we shall see, the functions and play a central role in the proofs of 8.1 and Theorems 4.1 and
From this point onward, we denote the true value of the tail index by
Theorem 4.1.
(existence and consistency) Assume that tail distribution satisfies the first-order condition and let be a continuous function fulfilling assumptions For a given sequence of integers such that and as with probability tending to there exists solution for the estimating equation such that as provided that
Since weak approximations of extreme value theory based statistics are achieved in the second-order framework (see e.g. de Haan and Stadtmüller, 1996), then it seems quite natural to suppose that satisfies the well-known second-order condition of regular variation. That is, we assume that for any
where denotes the second-order parameter, is a function tending to zero and not changing signs near infinity with regularly varying absolute value with index For convenience, we set and where for stands for the quantile function pertaining to cdf
Theorem 4.2.
(asymptotic normality) Assume that tail distribution satisfies the second-order condition and that, for the estimator is a consistent for satisfying Let be a continuous function fulfilling assumptions and be a sequence of integers such that and is asymptotically bounded. Then, in the probability space there exists a standard Wiener process such that
where and are respectively given in and and
Whenever
as where
5. Simulation study
In this section, we address the following two points:
The comparison of these estimators is conducted under the following two distributional models:
Burr distribution: defined by for and
Fréchet distribution: defined by for and
To investigate the first issue, we use the triweight function defined in and set the Burr distribution parameters to and We consider four values for the robustness parameter , namely and generate samples of size Figures 9.1 and 9.2 illustrate the behavior of the estimators and as a function of the number Further, Figures 9.3 and 9.6 display the absolute bias (ABIAS) and mean squared error (MSE) of and for values and with values of and These results are based on samples of size and also presented as a function of The simulation outcomes, as depicted in Figures 9.1 to 9.6, demonstrate that the estimator consistently outperforms particularly for larger values of While estimator becomes increasingly sensitive as the maintains greater stability, particularly as the true value of Furthermore, both ABIAS and MSE increase as the number of extreme values increases.
We now turn to the second issue by considering an -contaminated model as a mixture of two cdf’s and where denotes the outliers contamination fraction. We examine four mixture scenarios from the following models:
Burr contaminated by Burr
Fréchet contaminated by Fréchet
Fréchet contaminated by Burr
Burr contaminated by Fréchet
We choose set , and consider three values for namely For each scenario, we generate samples of size We assess the estimators and in terms of ABIAS and MSE under both contaminated and non-contaminated distributions as a functions of the number
Regarding the four scenarios depicted in Figures 9.7-9.10, we observe that introducing contamination affects both estimators in all cases. However, the estimator still outperforms the estimator as indicated by the ABIAS and MSE values. Moreover, the performance of both estimators improves as the number increases, even in the presence of contamination.
6. Proofs
Let us begin to define the following expressions: for
and
denotes the values of -th derivative of with respect to at
6.1. Proof of Theorem 4.1
To show the existence and consistency of we will adapt the proof of Theorem 1in Dierchx et al. (2013) which itself is an adaptation of the proof of Theorem 5.1 in Chapter 6 of Lehmann and Casella (1998). That theorem establishes the existence and consistency of solutions to the likelihood equations, and we extend it here to the WMDPDE framework. Let denotes the empirical counterpart of the weighted density power divergence objective function given in namely
Next we show that
where for By applying Taylor’s expansion near to function we decompose into
where is between and In the first assertion of Lemma 8.2 we stated that as In other terms, for any sufficiently small we have which entails that for ever with probability tending to From the second assertion, we have as where is that given in therefore The constant being positive, then there exists and such that for such that Note that is a consistent estimator for then by Lemma 8.4 we deduce that This means that with probability tending to there exists a constant such that Therefore
with probability tending to The choice implies that which leads to inequality To complete the proof of the existence and consistency, we follow the same steps to those used in the proof of Theorem 3.7 in Chapter 6 of Lehmann and Casella (1998). Let be small so that and then consider the set
We already showed that for any such then there exists a sequence such that as Note that being differentiable on then given there exists a point for which attains a local minimum, thereby Let us set for and arbitrary otherwise. Obviously
thus with probability tending to there exists a sequence of solutions to estimating equation Observe now that for any fixed and sufficiently large as which establishes the consistency of as sought.
6.2. Proof of Theorem 4.2
Let us retain the notations used in the proof of Theorem 4.1, namely those of and Applying Taylor’s expansion to the estimating equation yields
where is between and We showed above that it follows that as thus in view of Lemma 8.4 we get This implies that thereby
as Once again making use of Gaussian approximation of Lemma yields
where is a standard Wiener process and being the asymptotic bias given in Theorem From the second assertion of Lemma we have therefore as where
which equals this completes the proof of Theorem
Conclusion.
In this work, we proposed a new class of robust tail index estimators for Pareto-type distributions by incorporating a weight function into the density power divergence framework. This approach provides a flexible and effective generalization of existing weighted least squares and kernel-based estimators, enhancing robustness to outliers and model deviations. We established the asymptotic properties of the proposed estimators, including consistency and asymptotic normality, under general regularity conditions. A comprehensive simulation study demonstrated the improved finite-sample performance of our estimators compared to classical methods, particularly in the presence of contamination or small departures from the assumed model. These results highlight the relevance of robust statistical techniques in extreme value theory and open promising directions for future research, including extensions to multivariate settings and dependent data structures.
7. Declarations
The authors have no relevant financial or non-financial interests to disclose.
The authors have no competing interests to declare that are relevant to the content of this article.
All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
The authors have no financial or proprietary interests in any material discussed in this article.